A light source with a frequency of 1.8 x 1015 s-1 is introduced to a metal which has a work function of 5.9 x 10-19 J.

How much energy does the exiting electron have? J (eg. 7.2E-4)

What part of the electromagnetic radiation does the energy belong to?

What is the wavelength of a neutron that is traveling at a velocity of 5.8 x 103 m/s ?
m (eg. 3.24E-3)

Determine the number of sigma and pi bonds in the following structure?







σ bonds


∏ bonds


How many sigma and pi bonds are in the following structure (lone pairs not shown):



σ bonds

∏ bonds
1
Expert's answer
2020-12-06T13:29:09-0500

Given:

Potential of the electrons, V = 30 kV = 3 × 104 V

Speed of light, c = 3 × 108 m/s

Energy of the electrons, E = 3 × 10eV

Where, e = Charge on an electron = 1.6 × 10−19 C

  • Maximum frequency produced by the X-rays = ν

The energy of the electrons is given by the relation:

E = hν

Where, h = Planck’s constant = 6.626 × 10−34 Js

Therefore v = (E/h)

= (3 × 104 x 1.6 × 10−19)/ (6.626 × 10−34)

=7.24 x 1018 Hz

Hence, the maximum frequency of X-rays produced is 7.24 x 1018 Hz

  • The minimum wavelength produced by the X-rays is given as:

λ = (c/ν)

= (3 x 108)/ (7.24 x 1018)

=4.14 x 10-11 m

λ =0.0414 nm

Hence, the minimum wavelength of X-rays produced is 0.0414 nm.